# Predicting the Dynamics of Complex System via Multiscale Diffusion Autoencoder

### v1

@@ Abstract.
Complex System, Multiscale, Diffusion Model, Graph Neural ODE
@@ 1. Introduction
The dynamics of complex systems emerge from the nonlinear interactions and co-evolution of numerous components, giving rise to intricate spatiotemporal patterns and multiscale structures, as seen in fluid flow (Kochkov et al., 2021), bioproteins (Bhatia et al., 2021), and brain neurons (Brooks et al., 2009). Predicting the dynamics of such systems is crucial for real-world applications, including policy-making, resource management, and strategic planning (Preisler and Westerling, 2007; Brown et al., 2015; Bevacqua et al., 2023). In previous studies, the remarkable complexity of these systems has shown potential for reduction through dimensionality reduction techniques (Gao, 2024). The latent variables that bridge high-dimensional system states are sufficient to describe the emergent insightful phenomena arising from complex microscopic dynamics (Thibeault et al., 2024). Guided by this idea, a core objective of complex system dynamics prediction is to discover the low-dimensional latent space where the system’s intrinsic dynamics reside (Vlachas et al., 2022; Kontolati et al., 2024; Wu et al., 2024b).
@@ 2.2. Diffusion Model
Diffusion model (Sohl-Dickstein et al., 2015; Dhariwal and Nichol, 2021; Rombach et al., 2022) perturbs the data distribution by adding noise and learn to reverse the process through denoising, demonstrating strong fitting capabilities for data distributions in images, videos, and general sequences (Yang et al., 2023; Cao et al., 2024; Croitoru et al., 2023). We denote the original sample as $\bm{x}_{0}$ and the sample after $n$ diffusion steps as $\bm{x}_{n}$. In a standard diffusion model, the forward diffusion process iteratively adds noise to perturb the data: $\bm{x}_{n}=\sqrt{\overline{\alpha}_{n}}\bm{x}_{0}+\sqrt{1-\overline{\alpha}_{n}}\bm{\epsilon}$, where $\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})$ and $\{\alpha_{n}\}$ are noise schedules (Ho et al., 2020). The reverse process starts from Gaussian noise and progressively denoising to sample the real data point as
@@ 4. Experiments
In this section, we present the extensive evaluation results of MDPNet. We analyze the prediction performance of all models on four representative complex systems, followed by additional experiments to analysis MDPNet’s hyperparameter sensitivity, robustness, interpretability, generalization, ablation study and computational cost.
@@ 4.2. Baselines
We categorize these baselines into two types based on the state space where the dynamics prediction is executed: observed-space and latent-space, as shown in Table 1. In latent-space methods, predictions occur in the latent space after a encoding or downsampling.
@@ 4.3. Setup
For all models, we split the dataset into an 8:2 training-to-validation ratio. During training, the loss is calculated based on the prediction results at 5-step intervals, while during testing, we autoregressively predict the entire trajectory starting from the initial conditions. We evaluate the predicted trajectory against the true trajectory using normalized mean squared error (NMSE) (Ohana et al., 2024) and structural similarity index (SSIM) (Hore and Ziou, 2010) to assess the error at each grid point and the global structural similarity. NMSE is sensitive to numerical anomalies, while SSIM captures the structural and textural patterns of spatiotemporal dynamics, making the two metrics complementary and effective for evaluating prediction quality (details in the Appendix A). MDPNet and G-LED use 1,000 diffusion steps, with the scale schedule $\{N_{k}\}^{K-1}_{k=1}$ for MDPNet being uniformly distributed. Unless otherwise specified, we report the results for MDPNet with scale numer $K=3$.
4.4. Prediction Evaluation
We report the evaluation results of all models across four systems in Table 1. Compared to suboptimal models, MDPNet achieves an average improvement of 53.23% in NMSE and 5.16% in SSIM. This demonstrates, on one hand, that MDPNet successfully captures the intrinsic dynamics of complex systems and reliably reconstructs the original scale. On the other hand, it also validates the importance of incorporating multiscale information and modeling its interactions for accurate dynamical predictions.
In more than half of the cases, latent-space methods exhibit greater stability than observed-space methods (particularly in the Bruss and NS systems). This suggests that encoding-based approaches have greater potential in discovering the latent space where complex system dynamics reside, enabling more reliable predictions. However, in certain scenarios, such as the LO system, encoding-based methods generally show degradation. This may be attributed to the limitations of vanilla autoencoders, which are trained based on reconstruction loss and struggle to encode the vast degrees of freedom in complex spatiotemporal patterns, thereby becoming a bottleneck for prediction accuracy. In contrast, MDPNet, leveraging its novel multiscale diffusion autoencoder, effectively captures a meaningful latent space across all systems. We further analyze the advantages of the multiscale diffusion autoencoder in Sec. 4.7 and the cross-scale neural dynamics module in Sec. 4.9.
@@ 5.2. Multiscale Modeling of Dynamics Prediction
Complex systems consist of numerous interacting components, such as molecular particles in a chemical reaction or neurons in the brain (Wang et al., 2024b). The nonlinear interactions and feedback mechanisms at the microscopic level give rise to emergent ordered structures at the macroscopic scale, motivating the exploration of self-organizing dynamical mechanisms (Haken, 2006, 2012) and cross-scale co-evolution from a multiscale perspective (Peng et al., 2021; Cash et al., 2006). Bhatia et al. (2021) performs adaptive multiscale simulations of the interaction between RAS proteins and the plasma membrane. The simulation employs dynamic density functional theory at the macroscopic level and molecular dynamics simulations at the microscopic level. Vlachas et al. (2022) uses an autoencoder to extract the macroscopic state of a high-dimensional PDE system and alternately predicts dynamics at both the macroscopic and microscopic scales. Li et al. (2023) differentiates the fast and slow components in the overall dynamics based on changes in intrinsic dimensions, and uses Koopman operators and autoregressive models to predict these components separately. These methods independently predict dynamics at individual scales at each time step, neglecting the propagation of information across scales. Wang et al. (2020) generates coarse-grained fluid dynamics data using a reduced-order model and then employs a neural network to model the correlation between the coarse data and the observed fine-scale data. MultiScaleGNN (Lino et al., 2021) utilizes a U-Net-based graph neural network to implicitly capture multiscale information and propagates messages through edge connections. Compared to these works, our model explicitly decouples multiscale representations while preserving cross-scale interactions.
@@ 5.3. Diffusion Models for Dynamics Prediction
The tremendous success of diffusion models in video generation (Tumanyan et al., 2023; Jin et al., 2024) and time series modeling (Fan et al., 2024; Shen et al., 2024) has inspired a series of works in complex system dynamics prediction. Shu et al. (2023) and Li et al. (2024b) have leveraged large-scale pre-trained diffusion models to reconstruct high-fidelity data from conventional low-fidelity samples or sparse measurement data. Known partial differential equations provide physical conditioning information for the denoising process, enhancing accuracy. Building upon state-of-the-art diffusion models, Li et al. (2024a) propose a machine learning approach to generate single-particle trajectory data in high Reynolds number three-dimensional turbulence. Similarly, Lienen et al. (2023) treats turbulence simulation as a generative task, using a diffusion model to capture the distribution of turbulence induced by unseen objects and generate high-quality samples for downstream applications. A recent study (Bastek et al., 2024) introduces a first-principles-based loss term as physical knowledge to enhance the training process of the diffusion model, thus generating data samples that satisfy physical constraints. Rühling Cachay et al. (2024) aligns the temporal axis of spatiotemporal dynamics with diffusion’s process, replacing the noise injection with temporal interpolation and denoising with prediction, thereby embedding dynamical information into the diffusion process. G-LED (Gao et al., 2024) simply subsamples the system states as prediction targets, and uses predicted future frames as conditions for the diffusion model to reconstruct the high-fidelity original states. In contrast to these works, we combine a scale residual encoder with diffusion to collaboratively discover the latent space of complex systems. This multiscale information gradually guides the denoising process of diffusion, thereby improving reconstruction quality.
@@ 6. Conclusions
In terms of limitations, our model adopts a UNet-based conditional diffusion model. Integrating transformer-based latent diffusion models (Peebles and Xie, 2023) could improve scalability, paving the way for large-scale pre-trained models in spatiotemporal prediction. Future work will explore advanced architectures to improve both efficiency and scalability.

### v2

@@ Abstract.
Complex System, Multiscale, Diffusion Model
journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737087isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Multiscale systemsccs: Applied computing Physicsccs: Computing methodologies Learning latent representations
KDD Availability Link: The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.15574207.
@@ 1. Introduction
The dynamics of complex systems emerge from the nonlinear interactions and co-evolution of numerous components, giving rise to intricate spatiotemporal patterns and multiscale structures, as seen in fluid flow (Kochkov et al., 2021), bioproteins (Bhatia et al., 2021), and brain neurons (Brooks et al., 2009). Predicting the dynamics of such systems is crucial for real-world applications, including policy-making, resource management, and strategic planning (Preisler and Westerling, 2007; Brown et al., 2015; Bevacqua et al., 2023). In previous studies, the remarkable complexity of these systems has shown potential for reduction through dimensionality reduction techniques (Gao, 2024; Li et al., 2025b). The latent variables that bridge high-dimensional system states are sufficient to describe the emergent insightful phenomena arising from complex microscopic dynamics (Thibeault et al., 2024). Guided by this idea, a core objective of complex system dynamics prediction is to discover the low-dimensional latent space where the system’s intrinsic dynamics reside (Vlachas et al., 2022; Kontolati et al., 2024; Wu et al., 2024b; Ding et al., 2025).
@@ 2.2. Diffusion Model
Diffusion model (Dhariwal and Nichol, 2021; Rombach et al., 2022) perturbs the data distribution by adding noise and learn to reverse the process through denoising, demonstrating strong fitting capabilities for data distributions in images, videos, and general sequences (Yang et al., 2023; Cao et al., 2024; Croitoru et al., 2023; Sheng et al., 2025). We denote the original sample as $\bm{x}_{0}$ and the sample after $n$ diffusion steps as $\bm{x}_{n}$. In a standard diffusion model, the forward diffusion process iteratively adds noise to perturb the data: $\bm{x}_{n}=\sqrt{\overline{\alpha}_{n}}\bm{x}_{0}+\sqrt{1-\overline{\alpha}_{n}}\bm{\epsilon}$, where $\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})$ and $\{\alpha_{n}\}$ are noise schedules (Ho et al., 2020). The reverse process starts from Gaussian noise and progressively denoising to sample the real data point as
@@ 4. Experiments
In this section, we present the extensive evaluation results of MDPNet. We analyze the prediction performance of all models on four representative complex systems, followed by additional experiments to analyze MDPNet’s hyperparameter sensitivity, robustness, interpretability, generalization, ablation study, and computational cost.
@@ 4.2. Baselines
We categorize these baselines into two types based on the state space where the dynamics prediction is executed: observed-space and latent-space, as shown in Table 1. In latent-space methods, predictions occur in the latent space after encoding or downsampling.
@@ 4.3. Setup
For all models, we split the dataset into an 8:2 training-to-validation ratio. During training, the loss is calculated based on the prediction results at 5-step intervals, while during testing, we autoregressively predict the entire trajectory starting from the initial conditions. We evaluate the predicted trajectory against the true trajectory using normalized mean squared error (NMSE) (Ohana et al., 2024) and structural similarity index (SSIM) (Hore and Ziou, 2010) to assess the error at each grid point and the global structural similarity. NMSE is sensitive to numerical anomalies, while SSIM captures the structural and textural patterns of spatiotemporal dynamics, making the two metrics complementary and effective for evaluating prediction quality (details in the Appendix A). MDPNet and G-LED use 1,000 diffusion steps, with the scale schedule $\{N_{k}\}^{K-1}_{k=1}$ for MDPNet being uniformly distributed. Unless otherw
In contrast, MDPNet, leveraging its novel multiscale diffusion autoencoder, effectively captures a meaningful latent space across all systems. We further analyze the advantages of the multiscale diffusion autoencoder in Sec. 4.6 and the cross-scale neural dynamics module in Sec. 4.8.
@@ 5.2. Multiscale Modeling of Dynamics Prediction
Complex systems consist of numerous interacting components, such as molecular particles in a chemical reaction or neurons in the brain (Wang et al., 2024b). The nonlinear interactions and feedback mechanisms at the microscopic level give rise to emergent ordered structures at the macroscopic scale, motivating the exploration of self-organizing dynamical mechanisms (Haken, 2006, 2012) and cross-scale co-evolution from a multiscale perspective (Peng et al., 2021; Cash et al., 2006). Bhatia et al. (2021) performs adaptive multiscale simulations of the interaction between RAS proteins and the plasma membrane. The simulation employs dynamic density functional theory at the macroscopic level and molecular dynamics simulations at the microscopic level. Vlachas et al. (2022) uses an autoencoder to extract the macroscopic state of a high-dimensional PDE system and alternately predicts dynamics at both the macroscopic and microscopic scales. Li et al. (2023) differentiates the fast and slow components in the overall dynamics based on changes in intrinsic dimensions, and uses Koopman operators and autoregressive models to predict these components separately. These methods independently predict dynamics at individual scales at each time step, neglecting the propagation of information across scales. Wang et al. (2020) generates coarse-grained fluid dynamics data using a reduced-order model and then employs a neural network to model the correlation between the coarse data and the observed fine-scale data. MultiScaleGNN (Lino et al., 2021) utilizes a U-Net-based graph neural network to implicitly capture multiscale information and propagates messages through edge connections. In addition, numerous studies (Gupta et al., 2022; Xiao et al., 2022; Gupta et al., 2021) leverage wavelet transforms for the analysis of multi-scale features. Compared to these works, our model explicitly decouples multiscale representations while preserving cross-scale interactions.
@@ 5.3. Diffusion Models for Dynamics Prediction
The tremendous success of diffusion models in video generation (Tumanyan et al., 2023; Jin et al., 2024) and time series modeling (Fan et al., 2024; Shen et al., 2024) has inspired a series of works in complex system dynamics prediction (Li et al., 2025a). Shu et al. (2023) and Li et al. (2024b) have leveraged large-scale pre-trained diffusion models to reconstruct high-fidelity data from conventional low-fidelity samples or sparse measurement data. Known partial differential equations provide physical conditioning information for the denoising process, enhancing accuracy. Building upon state-of-the-art diffusion models, Li et al. (2024a) propose a machine learning approach to generate single-particle trajectory data in high Reynolds number three-dimensional turbulence. Similarly, Lienen et al. (2023) treats turbulence simulation as a generative task, using a diffusion model to capture the distribution of turbulence induced by unseen objects and generate high-quality samples for downstream applications. A recent study (Bastek et al., 2024) introduces a first-principles-based loss term as physical knowledge to enhance the training process of the diffusion model, thus generating data samples that satisfy physical constraints. Rühling Cachay et al. (2024) aligns the temporal axis of spatiotemporal dynamics with diffusion’s process, replacing the noise injection with temporal interpolation and denoising with prediction, thereby embedding dynamical information into the diffusion process. G-LED (Gao et al., 2024) simply subsamples the system states as prediction targets, and uses predicted future frames as conditions for the diffusion model to reconstruct the high-fidelity original states. In contrast to these works, we combine a scale residual encoder with diffusion to collaboratively discover the latent space of complex systems. This multiscale information gradually guides the denoising process of diffusion, thereby improving reconstruction quality.
@@ 6. Conclusions
In terms of limitations, this paper focuses on grid-based data. A direction for future work is to extend MDPNet’s multiscale concept to spatiotemporal dynamics prediction on graph-structured data.
Acknowledgements.
This work is supported in part by the National Key Research and Development Program of China under 2024YFC3307603, and the National Natural Science Foundation of China under 62171260.
@@ B.2. Residual Snapshots
Table 2. Prediction performance on the Swift–Hohenberg equation.
NMSE×$\times 10^{-2}$ SSIM×$\times 10^{-1}$
ConvLSTM 5.214 8.090
G-LED 3.750 9.087
Ours 2.043 9.477
B.3. Ablation study on the diffusion step
The original method uses uniform diffusion steps. We compare two heuristics: one based on the ratio of downsampled areas across scales, allocating steps from small to large (1:4:9) and from large to small (9:4:1). With 1,000 diffusion steps and 3 scales, uniform distribution (1:1:1) performs best (Table 3), ensuring each scale gets sufficient attention.
Table 3. Different diffusion processes.
NMSE×$\times 10^{-2}$ LO Bruss GS NS
1:4:9 2.444 0.179 2.841 1.454
9:4:1 2.682 0.244 3.227 1.241
1:1:1 1.145 0.044 1.144 1.154
Appendix C Model Architecture
Our architecture combines established components from generative modeling and graph learning:
C.1. Core Modules
  • Residual encoder $\phi_{k}$: A deep CNN with residual blocks and a linear output. Each block includes Conv, GroupNorm, SiLU, AvgPool, followed by a CBAM (Woo et al., 2018).
  • Diffusion decoder $\psi$: A UNet with conditional attention. Each layer contains multiple residual attention blocks. In the upsampling path, each residual attention block is followed by a conditional attention block, where $q$ is computed from the latent vector $z_{\tau_{k}}$, and $k$, $v$ are computed from the feature vector $x_{\tau,n_{k}}$.
  • Predictor-self $f_{\theta_{k}}$: A shared two-layer MLP with scale info in positional embeddings.
  • Predictor-inter $g_{\theta}$: A standard graph attention layer (Veličković et al., 2018).
C.2. Model Hyperparameters
  • Depth: $\phi_{k}$ has 2 residual blocks. $\psi$ has 3 layers, each with 2 residual attention blocks.
  • Width: Feature dimension for $\phi_{k}$ and predictors is 128. Diffusion decoder channels: [64, 128, 128].
C.3. Training Hyperparameters
We train the model for 1000 epochs using a learning rate of 1e-4 and a batch size of 18. The ODE solver is dopri5.